OpenStudy (anonymous):

Find the solution set of each of the following linear equations. a) 7x - 5y = 3 b) 3x_1 - 5x_2 + 4x_3 = 7

OpenStudy (anonymous):

its subscript.., just symbol for x

OpenStudy (anonymous):

\(x^2\) = x^2

OpenStudy (anonymous):

\(x_2 =\) = x_2

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

i told you that \(x^3 \) = x^3. I wrote x_3, and x_3 = \(x_3\)